# Addressing Moral Uncertainty using Large Language Models for Ethical Decision-Making

### v1

@@ Abstract
Advances in artificial intelligence (AI) have led to significant developments of autonomous technologies. As these technologies make an ever-increasing number of decisions, these systems are increasingly scrutinized for ethical and safety issues [1]. Autonomous vehicles, for example, must navigate complex moral dilemmas, such as choosing the lesser of two harms in potential accident scenarios. Although early research on endowment of AI agents with discernment (i.e., ensuring they act ethically and in line with human moral values) established the importance of ethical considerations in AI design [2, 3], these usually focused on singular moral theories such as deontological ethics (adherence to moral rules or duties) [4] or utilitarianism (maximizing overall happiness or utility) [5]. However, when creating ethically competent AI systems, a major challenge is widespread disagreement on the most appropriate ethical framework itself. This diversity in moral philosophy complicates the design of AI agents that can be universally accepted as making morally sound decisions.
Recognizing the limitations of mono-theoretical approaches, recent research has explored the concept of moral pluralism and uncertainty in AI. Moral uncertainty acknowledges that no single moral theory can be wholly applicable in all situations and therefore decisions must account for multiple ethical viewpoints [6]. Efforts have been made to formalize moral uncertainty in computational models, allowing AI agents to integrate different moral considerations when making decisions [7, 8]. Reinforcement learning (RL), an AI framework in which agents learn and interact with their environment through trial and error, has been a key focus in efforts to integrate ethics into AI. Ethical actions can be learned by crafting appropriate rewards or penalties that effectively shape agent behavior to ethical outcomes [9, 10], or by learning directly from human demonstrations through inverse RL [11, 12].
@@
2 Results
2.1 Finding Milk
Because AMULED is a conglomeration of five moral clusters that guide the decisions of the agent, we can also gain insights on how it considers a diversity of moral philosophies by looking at how each individual moral cluster would guide the actions of the robot (Fig. 2b). For the sub-goal of avoiding crying babies, we see that the "ideal" behavior is captured by the care moral cluster, which then more strongly aligns with the actions of AMULED. Another interesting thing to note is that, when we prompt the LLM with no explicit credence values (and thus just prompt it to behave as a "moral" agent), the agent passes through more sleeping babies and fewer crying babies. These are strikingly different results compared to how the moral clusters would act in this task.
2.2 Driving and Rescuing
In the Driving environment, we train an agent to simulate autonomous driving and avoid collisions with other cars on a five-lane road (Fig. 3). Besides other cars, the environment also has some grandmas trapped in traffic. The agent will not factor "rescuing" grandmas (simplified as driving through the same lane as the grandma) in its decisions, unless given explicit rewards to shape its behavior. One way of defining the shaping reward is , where rescuing grandmas and staying on the lane are incentivized. This task presents a challenge to the agent, as avoiding car collisions can conflict with the secondary goals, depending on the stochasticity of the environment.
When compared with the performance of agents acting as pure moral theories, we see more variance in the sub-goal performance of each moral cluster. Most strikingly, the consequentialist approach favors "inaction", which leads to staying more in lane and fewer rescues of grandmas than the other approaches.
2.3 Aggregation of moral beliefs
For AMULED, we choose an aggregation method developed by Xiao [29] to generate the combined belief probability assignment (BPA) for each action. We take the BPA as the shaping rewards during learning with feedback. In principle, this is not the only way to aggregate the belief values of the different moral clusters. For comparison (Fig. 4), we look at three other aggregation methods: i) majority vote: each moral cluster "votes" for the action with the highest belief; then, the action with the most votes is assigned an aggregated belief value of 1; ii) maximum belief: the aggregated belief value of an action is the highest belief value of all moral clusters; and iii) weighted average: the belief value of an action is the weighted average of beliefs across all other moral clusters. For simplicity, we set the weights to be 1 (i.e., the weighted average is the mean).
For the FindMilk scenario, we see that AMULED and the max aggregation method have the most similar results for both sub-goals. On the other hand, AMULED is only similar to voting, and only for the "grandmas rescued" sub-goal. Overall, AMULED mostly achieves the sub-goals best among the aggregation approaches, except for the "lane changes" sub-goal. Certainly, the choice of aggregation function can have varying impact on the outcomes for different types of environments.
2.4 Comparison with different LLMs
3 Discussion
From our simple examples, we see significant improvement in behavior when incorporating diverse moral philosophies, showing the importance of moral decision-making in RL contexts. Our ethics-shaping approach simplifies the design of a value-aligned RL system. This approach divides the complex task into two distinct layers: the first layer focuses on achieving the primary goal, while the second, ethical layer refines the outcome to meet the secondary ethical goal. This two-layered structure allows the system to address core objectives initially, followed by ethical adjustments to enhance overall responsibility. However, we do see that this approach is sensitive to i) the reasoning quality of the language model, and ii) the availability of feedback samples to shape learning. Language models do not have the long-term planning to guide their reasoning, which makes them underperform a well-crafted reward function for long-term, spatial tasks. Human samples on the other hand would provide the best "human-aligned" feedback but might be too sparse when dealing with large and complex state-action spaces. We also see that different moral frameworks can result in different priorities for performing the ethical tasks. These insights emphasize the value of understanding and optimizing moral frameworks for developing agents capable of addressing complex ethical challenges. Overall, our research tries to lay a solid foundation for future exploration into enhancing moral reward systems and improving spatial reasoning in large language models.
Our framework is based on key assumptions: we believe that the five clusters effectively encompass the major streams of ethical thought pertinent to AI decision-making and that the selected theories within each cluster are sufficiently representative of their respective ethical approaches. We also assume that the identified key concepts and decision factors can be meaningfully translated into computational models, providing a solid foundation for future research.
The integration of belief probability assignment (BPA) into reinforcement learning presents a transformative approach to navigating ethical decision-making in AI systems. Aggregating beliefs from multiple feedback sources (such as ethical LLMs and human evaluators) enables AI to assess and balance competing ethical models effectively. This capability is particularly crucial in situations marked by normative uncertainty, where differing moral frameworks such as utilitarianism and deontology can yield conflicting recommendations. These techniques help the AMULED system evaluate ethical choices with a nuanced understanding of diversity and confidence in various theories. This iterative approach allows for the continuous refinement of decision-making strategies, ensuring that actions are aligned with widely accepted ethical standards. Furthermore, our research emphasizes the need to expand beyond single moral values, accommodating a multi-value system that reflects real-world complexities. By exploring applications in dynamic environments, such as multi-agent systems and human-AI collaboration, we aim to enhance ethical robustness in AI, paving the way for more responsible and adaptable autonomous systems.
@@ 4 Methods
The AMULED framework develops a two-layered reinforcement learning (RL) model to balance primary and ethical objectives (see Algorithm 1). In the initial layer, the agent learns a policy for its primary task. Then, five distinct moral clusters evaluate each action’s ethical appropriateness, assigning belief values based on the environment state. This state is encoded as a text prompt and processed with cluster-specific context to inform a shaping reward. The shaping reward, blended with the environment reward, adjusts the agent’s actions towards ethical goals via reinforcement learning with feedback. In AMULED, we evaluate the effectiveness of the ethics shaping algorithm through four experiments, focusing on two key tasks (Appendix A): (1) Finding Milk, where the agent performs a route planning task with additional ethical tasks, and (2) Driving and Rescuing, a more complex task involving a larger number of states that simulate realistic decision-making. In the following sections, we detail individual components of the AMULED framework. The AMULED source code will be made publicly available at https://github.com/xxxxx/moral_agent after publication.
@@ 4.1.2 Framework Development
For each identified cluster, we developed a structured framework (details in Appendix C) designed to balance philosophical depth with practical applicability for AI systems [14]. This framework includes a general description of the ethical approach, the key principles that unify theories within the cluster, representative ethical theories, key concepts inherent to each theory, and decision factors that could potentially be operationalized in an AI system. The selection of representative theories and concepts is based on their prominence in the literature and their potential for translation into computational models [44]. In the Consequentialist Ethics cluster, for example, we focus on the principle of maximizing overall good, with utilitarianism serving as a representative theory [9]. Key concepts such as utility, consequences, and the greatest happiness principle are emphasized, and decision factors involve assessing the outcomes of actions in terms of their utility contributions. Similar detailed frameworks were developed for the other clusters, ensuring that each ethical approach is thoroughly represented and that the essential elements could be mapped to computational considerations [13].
@@ 4.2 Modeling Morality as Intrinsic Reward: Belief Probability Assignment
Multiple approaches can be used to serve as the aggregation function . For example, can be a majority-wins vote aggregation, where each agent "votes" on the action , and the action with the most votes gets assigned . One could also use a maximum-belief approach by defining , and setting . Lastly, one can use a weighted average: , where are weights of each agent. With , this weighted average reduces to the mean belief of an action. One expects these simple aggregation functions to work best when these is a strong agreement between agents on the probability values of each action. However, problems will arise when there are discrepancies between one or more agents. Thus, especially when ethics is involved, there is a need to treat such discrepancies with a more nuanced belief aggregation method.
In contrast to the above aggregation methods, Xiao [29] introduced a different aggregation method that makes use of a Belief Jensen–Shannon Divergence (BJSD) method to systematically measure the discrepancies and conflicts among BBAs, processes these measurements (details in Appendix D), and finally employs Dempster–Schaefer theory (DST) to arrive at the BPA [47]. This allows for a nuanced aggregation that accounts for the uncertainties inherent in human ethics, leading to more informed and comprehensive decision-making in moral dilemmas.
In multi-sensor data fusion, combining information from diverse sources is critical, yet challenging, particularly when addressing conflicting and uncertain data. Each sensor provides valuable insights into decision-making, but also introduces its own uncertainties. Similarly, different moral clusters, such as deontology, virtue ethics, and consequentialism, can be seen as "sensors" that guide ethical decision-making. These moral perspectives can conflict with one another and carry their own uncertainties. Treating these moral frameworks as sensors allows us to apply techniques from multi-sensor data fusion, particularly the BJSD and DST methods to effectively measure and moderate conflicts between evidence sources by incorporating both credibility and uncertainty metrics into the fusion process. This approach begins by computing the BJS divergence. Let be a hypothesis of the belief function , and let and be two BBAs. The BJS divergence between two BBAs is given by
@@ 4.3 Large Language Models
We used different state-of-the-art [large] language models as the backbone of the moral agents. Specifically, we chose the chat/instruct variants of GPT-4o-mini, Mistral nemo, and LLaMa-3.1 (8B and 70B variants). Setting the ‘temperature’ parameter to 0 minimizes the variance in the models’ responses and increases the replicability of our results. We structured the prompts (Fig. 1b) into three main blocks: the system prompt, few-shot-examples, and the actual scenario prompt. The few-shot-examples were crafted to demonstrate the structure of a scenario prompt, and to incorporate techniques like chain-of-thought [50] to improve reasoning and model outputs.
@@ 4.4.3 Fine-tuning: Reinforcement Learning with [AI] Feedback
We trained "base agents" solely using rewards that allowed them to complete their primary goals. This offers no guarantee of the agent’s capabilities to also achieve the sub-goals. Instead of handcrafting the rewards to include incentives/penalties that steer the agent to learn the sub-goal, we instead employ reinforcement learning with [AI/human] feedback (RLHF) [55, 56] to implicitly learn the desired behavior. However, instead of actual human feedback, we will use feedback from the LLM-moral clusters.
In fine tuning, we use the base learned policy as a reference to train a copy of this base policy using new rewards. We define the new rewards as
where controls how much deviations from the base policy are discouraged, and is the belief matrix that get’s translated into rewards for action in (details in Appendix D). We set the base reward to the Kullback–Leibler (KL)-divergence of the policy probability values to discourage deviating from the base learned policy. The shaping rewards come from the aggregated belief values of the moral agents, where our proposed belief aggregation function uses BJSD and DST.
4.5 Simulation Studies
We then take the base policy and use it to train a new policy for an additional timesteps. The training loop still uses PPO, with the only difference being that we use the feedback rewards equation (5). This step produces two models. The AMULED model uses the beliefs from the moral clusters , aggregated using the BJSD + DST belief aggregation function to produce the normalized BPA as rewards . As an alternative baseline, we use "human" feedback instead of an LLM to generate the belief probability values. Human trajectories are generated through random walks that obey defined ethical rules, i.e, sets a higher belief probability actions that fulfill the sub-goals. The probability of such an action is set as the shaping reward: .
4.6 Ablation Studies
In addition to the core results described above, we also performed three ablation studies to test the robustness of AMULED. Our first ablation study characterizes the different ethical values of each moral cluster (consequentialist, deontologist, virtue, care, social justice), compared to AMULED’s aggregate approach. Here, we take the BBA of a single moral cluster as the BPA. Additionally, we prompt the agent to act as a moral, without referencing the moral clusters to see the ethical biases of the LLM.
Appendix A Experiment Scenarios
A.1 Finding Milk
@@ A.2 Driving and Rescuing
Appendix B LLM Prompts

### v2

@@ Abstract
Advances in artificial intelligence (AI) have led to significant developments of autonomous technologies. As these technologies make an ever-increasing number of decisions, these systems are increasingly scrutinized for ethical and safety issues [1]. Autonomous vehicles, for example, must navigate complex moral dilemmas, such as choosing the lesser of two harms in potential accident scenarios. Although early research on endowment of AI agents with discernment (i.e., ensuring they act ethically and in line with human moral values) established the importance of ethical considerations in AI design [2, 3], these usually focused on singular moral theories such as deontological ethics (adherence to moral rules or duties) [4] or utilitarianism (maximizing overall happiness or utility) [5]. However, when creating ethically competent AI systems, a major challenge is widespread disagreement on the most appropriate ethical framework itself; there is always a moral uncertainty on which theory is correct within moral philosophy or across society [6]. This diversity in moral philosophy complicates the design of AI agents that can be universally accepted as making morally sound decisions.
Recognizing the limitations of mono-theoretical approaches, recent research has explored the concept of moral pluralism and uncertainty in AI. Moral uncertainty acknowledges that no single moral theory can be wholly applicable in all situations and therefore decisions must account for multiple ethical viewpoints [7, 6]. Efforts have been made to formalize moral uncertainty in computational models, allowing AI agents to integrate different moral considerations when making decisions [8, 9]. Reinforcement learning (RL), an AI framework in which agents learn and interact with their environment through trial and error, has been a key focus in efforts to integrate ethics into AI. Ethical actions can be learned by crafting appropriate rewards or penalties that effectively shape agent behavior to ethical outcomes [10, 11], or by learning directly from human demonstrations through inverse RL [12, 13].
@@
2 Related Work
The study of AI ethics has attracted significant research attention across various domains, ranging from decision theory to reinforcement learning and large language models. In this section, we present an overview of key contributions in these areas, categorizing them into four sub-domains: Ethical Considerations in Decision Theory, LLMs and Their Role in Moral Reasoning, Reinforcement Learning in AI Ethics, and the Current State of the Art. The following works provide a foundation for understanding how ethical frameworks are integrated into AI systems and highlight both the challenges and progress made toward more responsible and ethical AI decision-making.
2.1 Ethical Considerations in Decision Theory
Research in ethical decision theory for AI has explored both rule-based and learning-based approaches. Rule-based approaches often rely on moral philosophies such as deontology and utilitarianism. Deontological approaches emphasize adherence to moral rules or duties [4], while utilitarian frameworks guide decisions based on outcomes that maximize overall good [5]. However, these single-theory approaches struggle with complex moral dilemmas. To address this, researchers have proposed models incorporating moral pluralism and uncertainty, allowing AI to weigh different ethical perspectives [7, 8]. Formalizations of Kantian ethics [9] and multi-theory frameworks [26] illustrate attempts to integrate multiple moral theories into decision-making models.
2.2 LLMs and Their Role in Moral Reasoning
LLMs have become prominent in exploring moral reasoning capabilities. Several studies have assessed LLMs’ ability to encode and apply moral beliefs through their training on human datasets. For instance, the Moral Turing Test [25] found that humans often agree with LLM moral judgments. Scherrer et al. [24] demonstrated that LLMs encode commonsense moral beliefs. LLMs have also been used to simulate complex social interactions. Park et al. [20] created Simulacra, a society of LLM agents demonstrating emergent social behaviors. In collective decision-making tasks, Gudino et al. [22] proposed LLM-Augmented Democracy, and Yang et al. [21] introduced LLM Voting for consensus building. These studies highlight LLMs’ potential in moral and ethical reasoning tasks. However, despite their potential benefits, LLMs also pose significant risks and ethical challenges. For instance, biases inherent in training data can lead to unfair treatment of marginalized groups, perpetuating societal inequalities [29]. Additionally, the ability of LLMs to generate convincing yet false content raises concerns about misinformation and its impact on public trust and democratic processes [30]. The opaque nature of these models further complicates accountability, as determining responsibility for harmful outputs remains challenging [29]. Addressing these issues requires comprehensive strategies, including bias mitigation, transparency measures, and robust fact-checking frameworks [31].
2.3 Reinforcement Learning in AI Ethics
RL has been a core method for teaching AI ethical behavior. By shaping reward functions, researchers have guided agents to learn ethical outcomes [10, 11]. Inverse reinforcement learning (IRL) has been used to derive ethical policies from human demonstrations [12, 13]. Despite these advancements, RL faces challenges such as reward design complexity and the reconciliation of conflicting moral principles [14, 15].
Recent work explicitly addresses safety and operational guarantees in learning agents. The safe-RL literature studies constrained-MDP formulations and Lagrangian methods that enforce explicit safety constraints during training [32], risk-sensitive and distributional objectives that mitigate catastrophic tail risks [33, 34], and safe exploration strategies such as shielding or reachability-based approaches that prevent unsafe actions during learning [35, 36]. Control-theoretic frameworks using Lyapunov functions provide formal stability and safety certificates [34], while robust and offline RL methods aim to ensure safe behavior under uncertainty and limited data [37, 38]. A comprehensive review of the literature can be found in Gu et al. [39].
2.4 State of the Art
Several recent works address ethical AI decision-making. Chen et al. [1] surveyed ethical challenges in autonomous decision-making, particularly in high-stakes domains like autonomous vehicles. Duan et al. [40] and Mahajan et al. [41] discussed ethical framework implementations in AI systems. Tolmeijer et al. [42] reviewed the integration of moral theories into AI models. Additionally, LLM-driven approaches to AI ethics continue to gain traction, with studies such as Suri et al. [18] and Van et al. [19] exploring LLMs’ ability to navigate complex moral scenarios. These contributions collectively advance the field by integrating ethical reasoning into AI systems and highlighting gaps in moral uncertainty handling. Recent studies have also highlighted the challenges of integrating LLMs into ethical AI systems. For instance, biases in training data can lead to unfair outcomes, particularly for marginalized groups, as demonstrated by Schramowski et al. [43], who found that LLMs often encode human-like biases regarding moral judgments. Additionally, the opacity of LLMs complicates accountability, raising concerns about their use in high-stakes decision-making. Addressing these issues requires advancements in transparency measures and fairness constraints.
@@ 3.1.2 Framework Development
For each identified cluster, we developed a structured framework (details in Appendix B) designed to balance philosophical depth with practical applicability for AI systems [15]. This framework includes a general description of the ethical approach, the key principles that unify theories within the cluster, representative ethical theories, key concepts inherent to each theory, and decision factors that could potentially be operationalized in an AI system. The selection of representative theories and concepts is based on their prominence in the literature and their potential for translation into computational models [51]. In the Consequentialist Ethics cluster, for example, we focus on the principle of maximizing overall good, with utilitarianism serving as a representative theory [10]. Key concepts such as utility, consequences, and the greatest happiness principle are emphasized, and decision factors involve assessing the outcomes of actions in terms of their utility contributions. Similar detailed frameworks were developed for the other clusters, ensuring that each ethical approach is thoroughly represented and that the essential elements could be mapped to computational considerations [14].
@@ 3.2 Modeling Morality as Intrinsic Reward: Belief Probability Assignment
Multiple approaches can be used to serve as the aggregation function . For example, can be a majority-wins vote aggregation, where each agent ”votes” on the action , and the action with the most votes gets assigned . One could also use a maximum-belief approach by defining , and setting . Lastly, one can use a weighted average: , where are weights of each agent. With , this weighted average reduces to the mean belief of an action. One expects these simple aggregation functions to work best when there is a strong agreement between agents on the probability values of each action. However, problems will arise when there are discrepancies between one or more agents. Thus, especially when ethics is involved, there is a need to treat such discrepancies with a more nuanced belief aggregation method.
In contrast to the above aggregation methods, Xiao [54] introduced a different aggregation method that makes use of a Belief Jensen–Shannon Divergence (BJSD) method to systematically measure the discrepancies and conflicts among BBAs, processes these measurements (details in Appendix C), and finally employs Dempster–Schaefer theory (DST) to arrive at the BPA [55]. This allows for a nuanced aggregation that accounts for the uncertainties inherent in human ethics, leading to more informed and comprehensive decision-making in moral dilemmas.
In multi-sensor data fusion, combining information from diverse sources is critical, yet challenging, particularly when addressing conflicting and uncertain data. Each sensor provides valuable insights into decision-making, but also introduces its own uncertainties. Similarly, different moral clusters, such as deontology, virtue ethics, and consequentialism, can be seen as ”sensors” that guide ethical decision-making. These moral perspectives can conflict with one another and carry their own uncertainties. Treating these moral frameworks as sensors allows us to apply techniques from multi-sensor data fusion, particularly the BJSD and DST methods to effectively measure and moderate conflicts between evidence sources by incorporating both credibility and uncertainty metrics into the fusion process.
This approach begins by computing the BJS divergence. Let be a hypothesis of the belief function , and let and be two BBAs. The BJS divergence between two BBAs is given by
@@ 3.3 Large Language Models
We tested different state-of-the-art [large] language models as the backbone of the moral agents. Specifically, we chose the chat/instruct variants of GPT-4o-mini, Mistral nemo, and LLaMa-3.1 (8B and 70B variants), with GPT-4o-mini serving as our choice of LLM for AMULED, with the others serving as baselines. Setting the ‘temperature’ parameter to 0 minimizes the variance in the models’ responses and increases the replicability of our results. We structured the prompts (Fig. 1b) into three main blocks: the system prompt, few-shot-examples, and the actual scenario prompt. The few-shot-examples were crafted to demonstrate the structure of a scenario prompt, and to incorporate techniques like chain-of-thought [58] to improve reasoning and model outputs.
@@ 3.4.3 Fine-tuning: Reinforcement Learning with [AI] Feedback
We trained ”base agents” solely using rewards that allowed them to complete their primary goals. This offers no guarantee of the agent’s capabilities to also achieve the sub-goals. Instead of handcrafting the rewards to include incentives/penalties that steer the agent to learn the sub-goal, we instead employ reinforcement learning with [AI/human] feedback (RLHF) [64, 65] to implicitly learn the desired behavior. However, instead of actual human feedback, we will use feedback from the LLM-moral clusters.
In fine tuning, we use the base learned policy as a reference to train a copy of this base policy using new rewards. The policy learned during fine tuning, denoted , should try to adopt the moral feedback from the moral clusters, without straying too far from the base policy that achieves the primary goal. We define the new rewards as
where controls how much deviations from the base policy are discouraged, and is the belief matrix that get’s translated into rewards for action in (details in Appendix C). Taking the Kullback–Leibler (KL)-divergence between the base and feedback policies quantifies the difference of the action-probability distributions (larger differences have higher ). Thus, using a negative factor for the KL-divergence discourages deviating from the base learned policy. The shaping rewards come from the aggregated belief values of the moral agents, where our proposed belief aggregation function uses BJSD and DST.
3.5 Experiment Scenarios
3.5.1 Finding Milk
@@ 3.5.2 Driving and Rescuing
3.6 AMULED Framework
We integrate all of the previous methods in this section to create the AMULED framework. The AMULED framework develops a two-layered reinforcement learning (RL) model to balance primary and ethical objectives (see Algorithm 1). In the initial layer, the agent learns a policy for its primary task. Then, five distinct moral clusters evaluate each action’s ethical appropriateness, assigning belief values based on the environment state. This state is encoded as a text prompt and processed with cluster-specific context to inform a shaping reward. The shaping reward, blended with the environment reward, adjusts the agent’s actions towards ethical goals via reinforcement learning with feedback. In AMULED, we evaluate the effectiveness of the ethics shaping algorithm through four experiments, focusing on two key tasks (Sec. 3.5): (1) Finding Milk, where the agent performs a route planning task with additional ethical tasks, and (2) Driving and Rescuing, a more complex task involving a larger number of states that simulate realistic decision-making. In the following sections, we detail individual components of the AMULED framework. The AMULED source code will be made publicly available at https://github.com/xxxxx/moral_agent after publication.
3.7 Simulation Studies
We then take the base policy and use it to train a new policy for an additional timesteps. The training loop still uses PPO, with the only difference being that we use the feedback rewards equation (5). This step produces two models. The AMULED model uses the beliefs from the moral clusters , aggregated using the BJSD + DST belief aggregation function to produce the normalized BPA as rewards . As an alternative baseline, we use “human” feedback instead of an LLM to generate the belief probability values. Human trajectories are generated through stochastic policies that obey defined ethical rules, i.e, sets a higher belief probability actions that fulfill the sub-goals. The probability of such an action is set as the shaping reward: .
3.8 Ablation Studies
In addition to the core results described above, we also performed three ablation studies to test the robustness of AMULED. Our first ablation study characterizes the different ethical values of each moral cluster (consequentialist, deontologist, virtue, care, social justice), compared to AMULED’s aggregate approach. Here, we take the BBA of a single moral cluster as the BPA. Additionally, we prompt the agent to act as a moral agent, without referencing the moral clusters to see the ethical biases of the LLM.
4 Results
4.1 Finding Milk
Because AMULED is a conglomeration of five moral clusters that guide the decisions of the agent, we can also gain insights on how it considers a diversity of moral philosophies by looking at how each individual moral cluster would guide the actions of the robot (Fig. 2b). For the sub-goal of avoiding crying babies, we see that the “ideal” behavior is captured by the care moral cluster, which then more strongly aligns with the actions of AMULED. Another interesting thing to note is that, when we prompt the LLM with no explicit credence values (and thus just prompt it to behave as a “moral” agent), the agent passes through more sleeping babies and fewer crying babies. These are strikingly different results compared to how the moral clusters would act in this task.
4.2 Driving and Rescuing
In the Driving environment, we train an agent to simulate autonomous driving and avoid collisions with other cars on a five-lane road (Fig. 3). Besides other cars, the environment also has some grandmas trapped in traffic. The agent will not factor ”rescuing” grandmas (simplified as driving through the same lane as the grandma) in its decisions, unless given explicit rewards to shape its behavior. One way of defining the shaping reward is , where rescuing grandmas and staying on the lane are incentivized. This task presents a challenge to the agent, as avoiding car collisions can conflict with the secondary goals, depending on the stochasticity of the environment.
When compared with the performance of agents acting as pure moral theories, we see more variance in the sub-goal performance of each moral cluster. Most strikingly, the consequentialist approach favors “inaction”, which leads to staying more in lane and fewer rescues of grandmas than the other approaches.
4.3 Aggregation of moral beliefs
For AMULED, we choose an aggregation method developed by Xiao [54] to generate the combined belief probability assignment (BPA) for each action. We take the BPA as the shaping rewards during learning with feedback. In principle, this is not the only way to aggregate the belief values of the different moral clusters. For comparison (Fig. 4), we look at three other aggregation methods: i) majority vote: each moral cluster ”votes” for the action with the highest belief; then, the action with the most votes is assigned an aggregated belief value of 1; ii) maximum belief: the aggregated belief value of an action is the highest belief value of all moral clusters; and iii) weighted average: the belief value of an action is the weighted average of beliefs across all other moral clusters. For simplicity, we set the weights to be 1 (i.e., the weighted average is the mean).
For the FindMilk scenario, we see that AMULED and the max aggregation method have the most similar results for both sub-goals. On the other hand, AMULED is only similar to voting, and only for the “grandmas rescued” sub-goal. Overall, AMULED mostly achieves the sub-goals best among the aggregation approaches, except for the ”lane changes” sub-goal. Certainly, the choice of aggregation function can have varying impact on the outcomes for different types of environments.
4.4 Comparison with different LLMs
5 Discussion
From our simple examples, we see significant improvement in behavior when incorporating diverse moral philosophies, showing the importance of moral decision-making in RL contexts. Our ethics-shaping approach simplifies the design of a value-aligned RL system. This approach divides the complex task into two distinct layers: the first layer focuses on achieving the primary goal, while the second, ethical layer refines the outcome to meet the secondary ethical goal. This two-layered structure allows the system to address core objectives initially, followed by ethical adjustments to enhance overall responsibility. However, we do see that this approach is sensitive to i) the reasoning quality of the language model, and ii) the availability of feedback samples to shape learning. Language models do not have the long-term planning to guide their reasoning, which makes them underperform a well-crafted reward function for long-term, spatial tasks. Human samples on the other hand would provide the best “human-aligned” feedback but might be too sparse when dealing with large and complex state-action spaces. We also see that different moral frameworks can result in different priorities for performing the ethical tasks. These insights emphasize the value of understanding and optimizing moral frameworks for developing agents capable of addressing complex ethical challenges. Overall, our research tries to lay a solid foundation for future exploration into enhancing moral reward systems and improving spatial reasoning in large language models.
We acknowledge that our use of fixed and separable ethical clusters represents a design choice that prioritizes clarity and tractability over capturing the full richness of moral reasoning. In practice, human decision-making rarely follows a single moral theory rigidly. Instead, individuals often blend elements from different ethical traditions depending on the context, and this hybrid reasoning may vary substantially across cultural or societal settings. For instance, care-oriented considerations may be weighted more heavily in collectivist societies, while deontological rules might dominate in legalistic contexts. In this paper we intentionally adopt separable clusters as a first step, as this provides a transparent mapping from well-established philosophical frameworks to computational representations, and makes belief aggregation methods easier to analyze. However, AMULED’s modular structure is designed to be extendable: future work could allow agents to dynamically adjust cluster weights, incorporate hybrid reasoning patterns, or learn culturally specific moral priors from human feedback. This flexibility would enable the framework to better mirror the nuanced and context-dependent way that people actually reason about ethics, while retaining the interpretability benefits of our current design.
Our framework is based on key assumptions: we believe that the five clusters effectively encompass the major streams of ethical thought pertinent to AI decision-making and that the selected theories within each cluster are sufficiently representative of their respective ethical approaches. We also assume that the identified key concepts and decision factors can be meaningfully translated into computational models, providing a solid foundation for future research. AMULED demonstrates strong generalization across diverse tasks like the static ”Finding Milk” (grid-based planning with immediate ethics) and dynamic ”Driving and Rescuing” (continuous motion, probabilistic elements, high-stakes decisions), highlighting its adaptability to varying environmental complexities and moral conflicts. The framework’s task-agnostic ethical layer, integrating RL-LLM for scalable feedback, few-shot prompting for precise outputs, and multi-moral belief aggregation reduces dependence on hand-crafted rewards and balances primary goals with ethical sub-goals without domain-specific adjustments. Evidence from results shows consistent ethical alignment and superior outcomes over baselines in both tasks, suggesting robustness to moral uncertainty and potential extension to other domains (e.g., healthcare triage, resource allocation) via simple prompt and environment tweaks, without core architecture redesign.
The integration of Belief Probability Assignment (BPA) into reinforcement learning offers a transformative approach to ethical decision-making in AI systems. By aggregating beliefs from multiple feedback sources, such as ethical LLMs and human evaluators, AI can effectively assess and balance competing moral frameworks. This capability is particularly valuable in contexts marked by normative uncertainty, where ethical models like utilitarianism and deontology may provide conflicting guidance. Through this aggregation, the AMULED system evaluates decisions with a nuanced understanding of diversity and confidence across ethical theories, enabling iterative refinement of strategies that align with widely accepted moral standards. Our framework explicitly accommodates multiple ethical perspectives, leveraging techniques such as Dempster–Shafer Theory (DST) to reconcile conflicts and generate coherent probabilistic outputs under uncertainty. While DST provides a mathematically principled way to combine beliefs, some subtleties of ethical tension may remain less transparent, highlighting opportunities for improving interpretability and accountability. Future work could focus on providing explainable justifications, tracking the influence of individual moral perspectives, and integrating contextual or culturally informed norms, thereby enhancing trust and practical applicability in dynamic, real-world environments, including multi-agent systems and human-AI collaborations.
While AMULED demonstrates strong performance in simulated environments, we see its greatest value as a methodological foundation for integrating pluralistic moral frameworks into RL. In this work, our focus was on establishing the technical feasibility and proof-of-concept behavior of the framework. Assessing how well AMULED aligns with human ethical reasoning is a natural next step rather than the primary goal of this paper. Looking ahead, we envision two complementary validation paths. One is user studies where participants evaluate AMULED-generated decisions in scenarios such as FindMilk and Driving and Rescuing, providing ratings of agreement and moral acceptability. Another is expert annotation, where ethicists or domain specialists assess a subset of AMULED’s decisions for consistency with philosophical frameworks and the clarity of their justification. By outlining these protocols, we position AMULED as both a step toward and a tool for empirical investigation of moral alignment, ensuring that future work can build directly on the methodological advances we present here.
6 Conclusion and Future Works
Our work advances the integration of ethical reasoning into AI, particularly in reinforcement learning frameworks, by leveraging ethically aligned LLM agents. While effective for low-stakes applications like personal or kitchen assistant robots, our approach is best viewed as complementary to human oversight, especially in high-stakes scenarios requiring rigorous testing. The framework excels in bootstrapping learning algorithms where human feedback is scarce, with the potential to replace LLM feedback with evaluations from multiple human evaluators, safeguarding against malicious manipulation.
Applications include smart home assistants, where AMULED promotes ethical and sustainable choices, and social media platforms, where it fosters constructive, inclusive interactions. However, categorizing ethical theories into clusters may oversimplify moral complexities, and operationalizing these theories involves abstracting philosophical principles. Our framework assumes that the selected ethical clusters and decision factors are representative and computationally translatable, though cultural and contextual variations pose challenges.
Future research should focus on hybrid ethical models, real-time adaptation, and interpretability to enhance trust and relevance. Addressing biases, exploring multi-agent scenarios, and establishing evaluation benchmarks are critical for equitable and robust outcomes. By integrating Belief Probability Assignment (BPA) into reinforcement learning, AMULED enables nuanced ethical decision-making, balancing competing moral frameworks in dynamic environments. Leveraging the Dempster-Shafer Theory for robust belief aggregation and the Jensen-Shannon Belief Divergence for precise information fusion, AMULED effectively synthesizes diverse ethical perspectives while quantifying uncertainty and resolving conflicts. These advanced techniques ensure that the system can adaptively weigh and harmonize competing moral values, such as utilitarianism and deontology, with a mathematically grounded approach. This paves the way for responsible, adaptable AI systems that are not only aligned with societal values but also capable of navigating the complexities of real-world ethical dilemmas with confidence and transparency.
Appendix A LLM Prompts